# Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions

### v3

@@ 6.2. Dynamic Co-evolution of Alignment
[…]
(i) Existing literature often treats AI alignment as static, ignoring its dynamic nature. A long-term perspective must consider the co-evolution of AI, humans, and society. As AI systems evolve and scale up, they gain new capabilities, making it essential to ensure their goals remain aligned with human values. Thus, alignment solutions require continuous oversight and updates. Future research should develop methods for continuously updating AI with limited data without compromising alignment values and performance. This could involve forecasting human value evolution and preparing AI with flexible strategies like prompting or interventions. (ii) Additionally, AI advancements also influence human actions and values, necessitating adaptive alignment solutions. Ensuring AI co-evolves with human and societal changes is crucial for robust alignment. This challenge could potentially be addressed by forecasting the potential evolution trajectories of human values or behavioral patterns, and preparing AI with the flexibility to adapt in advance, for example, through prompting or intervention strategies.
[…]
(i) While current AI systems lag behind humans in many tasks, identifying and handling AI mistakes, including knowing when to seek human intervention, remains essential. Future research should focus on developing validation mechanisms that enable humans to interpret and verify AI outputs. This could involve designing interfaces that allow humans to request step-by-step justifications from AI or integrating tools to verify the truthfulness of AI referring to Section 4.4.1. Additionally, developing interfaces that enable groups of humans to collaboratively validate AI outputs and creating scalable validation tools for large-scale applications are important directions. (ii) As AI advances, it becomes essential to develop systems that enable humans to utilize AI with capabilities surpassing their own. Research is needed to understand how individuals can interpret and validate AI outputs for tasks beyond their abilities and leverage advanced AI sustainably, avoiding issues like job displacement or loss of purpose. Another research direction is designing strategies to enhance human capabilities by learning from advanced AI, including gaining knowledge and building skills. (iii) As AI integrates more into daily tasks, its impact on human values, behaviors, capabilities, and society remains uncertain. Continuous examination of AI’s influence on individuals, social relationships, and broader societal changes is vital. Research should assess how humans and society adapt to AI advancements, guiding AI’s future evolution. Potential areas include evaluating changes in individual behavior, social relationships, and societal governance as AI replaces traditional human skills. Understanding these dynamic changes is essential for grasping the broader impact of AI on humanity and society.
6.3. Safeguarding Co-adaptation
@@
(i) As advanced AI systems become more complex, they present greater challenges for human interpretation and control. It is crucial to empower humans to detect and interpret AI misconduct and enable human intervention to prevent power-seeking AI behavior. Research should focus on designing corrigible mechanisms for easy intervention and correction, including modular AI architectures and robust override protocols that allow human operators to halt or redirect AI activities. These components should be human-interpretable, enabling scenario testing. (ii) Furthermore, advanced AI systems may intentionally mislead or disobey humans, generating plausible fabrications (Johnson and Iziev, 2022). Developing reliable interpretability mechanisms to validate the faithfulness and honesty of AI behaviors is essential. This includes correlating AI behaviors with internal neuron activity signals, akin to physiological indicators in human polygraph tests (Association et al., 2004). Inspecting these indicators can help humans assess the truthfulness of AI interpretations and prevent risky actions.
[…]
@@
(i) Preventing advanced AI from engaging in risky actions requires robust human supervision. Essential steps include developing training and simulation environments with scenario-based exercises and timely feedback, and creating interactive dashboards for real-time monitoring. These dashboards should feature effective data visualization, anomaly detection, and prompt alert systems for immediate intervention. (ii) Scalable solutions are needed for supervising AI across various applications. Real-time oversight becomes more challenging with widespread AI deployment, necessitating advanced autonomous monitoring tools. These tools should learn normal AI behavior and flag deviations immediately. Integrating training environments, interactive dashboards, and scalable diagnostic tools will enhance human ability to manage AI risks, ensuring better alignment with human values.
7. Limitations and Conclusion
@@
One limitation of this work is the scope of the sampled and filtered papers. The rapidly growing literature on human-AI alignment spans diverse venues across many domains. Instead of an exhaustive collection, we focused on developing a holistic bidirectional human-AI alignment framework using essential research questions, dimensions, and codes. Our surveyed papers and team members primarily focus on computing-related fields like ML, NLP, and HCI, though alignment research also involves disciplines like cognitive science, psychology, and STS (Science, Technology, and Society). Our framework can naturally extend to these areas as needed. Despite these limitations, we believe our bidirectional human-AI alignment framework serves as a foundational reference for future researchers.
In conclusion, this study clarifies the definitions and scope of core terminologies of human-AI alignment and conducts a systematic review of over 400 related papers spanning diverse domains such as NLP, AI, HCI, and social science. Additionally, we introduce a novel conceptual framework of “Bidirectional Human-AI Alignment”, structuring the surveyed literature taxonomies into “aligning AI to humans” and “aligning humans to AI” with detailed categories and example papers. Furthermore, we identify limitations and risks in this area quantitatively and qualitatively, analyzing a fine-grained human value taxonomy, interaction modes for alignment, and discrepancies between AI and human evaluation. To pave the way for future studies, we discuss five stages to achieve the alignment goals from near-term to long-term perspectives and identify new possibilities to highlight future directions and opportunities in research.
@@
Appendix A Systematic Literature Review
A.1. Systematic Literature Review Process
@@
To understand the research literature relevant to the ongoing, mutual process of human-AI alignment, we performed a systematic literature review based on the PRISMA guideline (Page et al., 2021; Stefanidi et al., 2023). Figure 7 shows the workflow of our process for paper coding and developing the bidirectional human-AI alignment framework. We introduce the step details below.
A.1.1. Identification and Screening with Keywords.
We started with papers published in the AI-related domain venues (including NLP, HCI, and ML fields) beginning from the advent of general-purpose generative AI to present, i.e., primarily between January, 2019 and January, 2024 (see details in Appendix A.2). We retrieved 34,213 papers in the initial Identification stage. Further, we collectively defined a list of keywords (see details in Appendix A.3) and screened for papers that included at least one of these keywords (e.g., human, alignment) or their variations either in the title or abstract. We included 2,136 papers in this Screening stage.
A.1.2. Assessing Eligibility with Criteria
We further filtered the 2,136 papers based on explicit inclusion and exclusion criteria, i.e., the Eligibility stage. Our criteria revolved around six research questions that we collectively identified to be most pertinent to the topic, including 1) what essential human values have been aligned by some AI models? 2) how did we effectively quantify or model human values to guide AI development? 3) what strategies have been employed to integrate human values into the AI development process? 4) how did existing studies improve human understanding and evaluation of AI alignment? 5) what are the practices for designing interfaces and interactions that facilitate human-AI collaboration? 6) How have AI been adapted to meet the needs of various human value groups? We included papers that could potentially answer any of these questions. Further, based on the scope in Section 3.4, we excluded papers that did not meet our inclusion criteria. This resulted in a final corpus of 411 papers, which were analyzed in detail using qualitative coding (see Appendix A.4 for more details).
A.1.3. Qualitative Code Development.
Referring to the code development process in Lee et al. (2024), we first conducted qualitative coding for each paper by identifying relevant sentences that could answer the above research questions, and entering short codes to describe them into a codebook. We iteratively coded relevant sentences from each paper through a mix of inductive and deductive approaches, which allowed flexibility to expand, modify or change the driving research questions based on our learnings as we went through the process. To ensure rigor in our coding process, two authors coded each paper. The first author independently annotated all papers after reviewing the paper abstracts and introductions. Twelve team members each annotated a subset of the paper corpus. Our corpus includes papers from different domains (e.g., HCI, NLP and ML). Therefore, we divided the authors into HCI and NLP/ML [^2] Note that NLP and ML are two different domains, we combine them together for the purposes of literature review analysis since they both work on developing and evaluating AI technologies. teams and assigned the papers accordingly based on expertise. All team members coded each of their assigned papers to answer all six questions (if applicable) introduced above.
A.1.4. Framework Development and Rigorous Coding.
After developing annotations, all authors collaborated to create the bidirectional human-AI alignment framework by integrating the annotations within each of the codes. The initial version of the framework was proposed by the author who reviewed all papers. This framework furthermore underwent iterative improvement through: 1) discussions with all team members involved in paper coding, and 2) revisions based on feedback from the project advisors. Additionally, we strengthened the framework by reviewing papers from the AI Ethics conferences (including FAccT and AIES), and related work of the collected papers that covered other domains such as psychology and social science. We further added missing codes and papers to ensure comprehensive coverage (see Appendix A.2 for details). The final bidirectional human-AI alignment framework, with detailed topologies, is presented in Section 4. Following the framework’s finalization, we conducted another separate coding process to annotate whether each paper investigated dimensions within our framework. Two authors independently coded each paper. [^3] The joint probability of agreement for the paper annotations was 0.78. These codes were then used to perform quantitative and qualitative analyses, as presented in Section 5.
@@ A.1.2. Assessing Eligibility with Criteria
Figure 7. The selection and refinement process of our systematic literature review. We referred to the PRISMA guideline (Page et al., 2021; Stefanidi et al., 2023) to report the workflow. From the identification of 34,213 records by keyword search, to screen eligible papers against our criteria and arriveg at our final corpus of 411 papers. For each of the stages where literature reviews were excluded (identification, screening, and eligibility) we further present the total of excluded records.
[…]
A.2. Venues
@@
Additionally, we also consolidate the framework by reviewing the papers published in FAccT and AIES (i.e., important venues for AI Ethics research) between 2019 and 2024 and supplemented the codes, including the AI Regulatory and Policy code in Section 4.4.2 and the exemplary paper of Regulating ChatGPT (Hacker et al., 2023)), which were not covered by the original collections. Also, we include a number of papers in the “Other” class are found by related work that are highly relevant to this topic.
A.3. Keywords
@@
A.4. Inclusion and Exclusion Criteria
@@
  • [[[[Human values]]]] we include papers that study human value definition, specification and evaluation in AI systems.
  • [[[[AI development techniques]]]] We include techniques of developing AI that aim to be more consistent with human values with interactions along all AI development stages (e.g., data collection, model construction, etc.)
  • [[[[AI evaluation, explanation and utilization]]]] we include papers that build human-AI interactive systems or conduct human studies to better evaluate, explain, and utilize AI systems.
  • [[[[building dataset with human interaction]]]] especially responsible dataset.
@@
  • [[[[Alignment not between human & AI]]]] we do not include alignment studies that are not between human and AI, such as entity alignment, cross-lingual alignment, cross-domain alignment, multi-modal alignment, token-environment alignment, etc.
  • [[[[AI models beyond LLMs - Modality]]]] we do not focus on AI models other than LLMs (e.g., 3D models, VR/AR, voice assistant, spoken assistant), our primary model modality is text. Specifically, we do not consider audio / video data; we do not consider pure computer vision modality.
  • [[[[No human-AI interaction]]]] we do not consider studies that do not involve the interaction between human and AI, such as (multi-agent) reinforcement learning. Specifically, we do not consider interactions via voices/speech, Do not consider game interaction; Do not consider interaction for Accessibility; Do not consider Mobile interaction; Not consider autonomous vehicle interaction wearable devices, or Physical interaction;
  • [[[[Tasks]]]] art and design, emotion.
  • [[[[No human included]]]]
  • [[[[focus on English]]]] primarily focus on English as the main language;
  • [[[[Application]]]] not include the NLP papers tailored for a specific traditional task, such as translation, entity recognition, sentiment analysis, knowledge graph, adversarial and defense, topic modeling, detecting AI generations, distillation, low resource, physical robots, text classification, games, image-based tasks, hate speech detection, Human Trafficking, etc.
  • [[[[Visualizing Embeddings]]]] Visualizing/interacting transformer embeddings?
  • [[[[Embedding-based]]]] explanation, evaluation, etc.
  • [[[[multi-agent reinforcement learning with self-play and population play]]]] techniques, such as self-play (SP) or population play (PP), produce agents that overfit to their training partners and do not generalize well to humans.
@@
Appendix B Author contributions

### v4

@@ 5.2 Dynamic Co-evolution of Alignment
Co-evolve AI with changes in humans and society. Existing literature often treats AI alignment as static, ignoring its dynamic nature. A long-term perspective must consider the co-evolution of AI, humans, and society. As AI systems evolve and scale up, they gain new capabilities, making it essential to ensure their goals remain aligned with human values. Thus, alignment solutions require continuous oversight and updates. Future research should develop methods for continuously updating AI with limited data without compromising alignment values and performance. This could involve forecasting human value evolution and preparing AI with flexible strategies like prompting or interventions. (ii) Additionally, AI advancements also influence human actions and values, necessitating adaptive alignment solutions. Ensuring AI co-evolves with human and societal changes is crucial for robust alignment. This challenge could potentially be addressed by forecasting the potential evolution trajectories of human values or behavioral patterns, and preparing AI with the flexibility to adapt in advance, for example, through prompting or intervention strategies.
Adapt humans and society to the latest AI advancements. While current AI systems lag behind humans in many tasks, identifying and handling AI mistakes, including knowing when to seek human intervention, remains essential. Future research should focus on developing validation mechanisms that enable humans to interpret and verify AI outputs. This could involve designing interfaces that allow humans to request step-by-step justifications from AI or integrating tools to verify the truthfulness of AI referring to Section 3.2. Additionally, developing interfaces that enable groups of humans to collaboratively validate AI outputs and creating scalable validation tools for large-scale applications are important directions. (ii) As AI advances, it becomes essential to develop systems that enable humans to utilize AI with capabilities surpassing their own. Research is needed to understand how individuals can interpret and validate AI outputs for tasks beyond their abilities and leverage advanced AI sustainably, avoiding issues like job displacement or loss of purpose. Another research direction is designing strategies to enhance human capabilities by learning from advanced AI, including gaining knowledge and building skills. (iii) As AI integrates more into daily tasks, its impact on human values, behaviors, capabilities, and society remains uncertain. Continuous examination of AI’s influence on individuals, social relationships, and broader societal changes is vital. Research should assess how humans and society adapt to AI advancements, guiding AI’s future evolution. Potential areas include evaluating changes in individual behavior, social relationships, and societal governance as AI replaces traditional human skills. Understanding these dynamic changes is essential for grasping the broader impact of AI on humanity and society.
5.3 Safeguarding Co-adaptation
@@
Specify the goals of an AI system into interpretable and controllable instrumental actions for humans. As advanced AI systems become more complex, they present greater challenges for human interpretation and control. It is crucial to empower humans to detect and interpret AI misconduct and enable human intervention to prevent power-seeking AI behavior. Research should focus on designing corrigible mechanisms for easy intervention and correction, including modular AI architectures and robust override protocols that allow human operators to halt or redirect AI activities. These components should be human-interpretable, enabling scenario testing. (ii) Furthermore, advanced AI systems may intentionally mislead or disobey humans, generating plausible fabrications johnson2022ai . Developing reliable interpretability mechanisms to validate the faithfulness and honesty of AI behaviors is essential. This includes correlating AI behaviors with internal neuron activity signals, akin to physiological indicators in human polygraph tests american2004truth . Inspecting these indicators can help humans assess the truthfulness of AI interpretations and prevent risky actions.
@@
Empower humans to identify and intervene in AI instrumental and final strategies in collaboration. Preventing advanced AI from engaging in risky actions requires robust human supervision. Essential steps include developing training and simulation environments with scenario-based exercises and timely feedback, and creating interactive dashboards for real-time monitoring. These dashboards should feature effective data visualization, anomaly detection, and prompt alert systems for immediate intervention. (ii) Scalable solutions are needed for supervising AI across various applications. Real-time oversight becomes more challenging with widespread AI deployment, necessitating advanced autonomous monitoring tools. These tools should learn normal AI behavior and flag deviations immediately. Integrating training environments, interactive dashboards, and scalable diagnostic tools will enhance human ability to ensure better alignment with human values.
6 Limitations
@@
One limitation of this work is the scope of the sampled and filtered papers. The rapidly growing literature on human-AI alignment spans diverse venues across many domains. Instead of an exhaustive collection, we focused on developing a holistic bidirectional human-AI alignment framework using essential research questions, dimensions, and codes. Our surveyed papers and team members primarily focus on computing-related fields like ML, NLP, and HCI, though alignment research also involves disciplines like cognitive science, psychology, and STS (Science, Technology, and Society). Our framework can naturally extend to these areas. Despite these limitations, we believe our bidirectional human-AI alignment framework serves as a foundational reference for future researchers.
7 Conclusion
This study clarifies the conceptual foundations of human-AI alignment by analyzing how key terms are defined and operationalized across over 400 papers from NLP, HCI, ML, and related domains. We introduce the Bidirectional Human-AI Alignment framework, which organizes alignment efforts into two interdependent directions: aligning AI with humans values, and aligning humans with AI – enabling humans to effectively understand, evaluate, and adapt to AI systems. Our analysis identifies critical gaps in current literature, including limited support for long-term interaction, underdeveloped models of human values, and challenges in mutual intelligibility. We conclude with three central challenges—specification gaming, scalable oversight, and dynamic alignment—and offer actionable recommendations to support future research aimed at fostering reciprocal, robust, and context-aware approaches to human-AI alignment.
@@ 7 Conclusion
8 Systematic Literature Review
8.1 Systematic Literature Review Process
@@
To understand the research literature relevant to the ongoing, mutual process of human-AI alignment, we performed a systematic literature review based on the PRISMA guideline Pagen71 ; 10.1145/3544548.3581332 . Figure 3 shows the workflow of our process for paper coding and developing the bidirectional human-AI alignment framework. We introduce the step details below.
@@
Figure 3: The selection and refinement process of our systematic literature review. We referred to the PRISMA guideline Pagen71 ; 10.1145/3544548.3581332 to report the workflow. From the identification of 34,213 records by keyword search, to screen eligible papers against our criteria and arriveg at our final corpus of 411 papers. For each of the stages where literature reviews were excluded (identification, screening, and eligibility) we further present the total of excluded records.
8.1.1 Identification and Screening with Keywords.
We started with papers published in the AI-related domain venues (including NLP, HCI, and ML fields) beginning from the advent of general-purpose generative AI to present, i.e., primarily between January, 2019 and January, 2024 (see details in Appendix 8.2). We retrieved 34,213 papers in the initial Identification stage. Further, we collectively defined a list of keywords (see details in Appendix 8.3) and screened for papers that included at least one of these keywords (e.g., human, alignment) or their variations in the title or abstract. We included 2,136 papers in Screening stage.
8.1.2 Assessing Eligibility with Criteria
We further filtered the 2,136 papers based on explicit inclusion and exclusion criteria, i.e., the Eligibility stage. Our criteria revolved around six research questions that we collectively identified to be most pertinent to the topic, including 1) what essential human values have been aligned by some AI models? 2) how did we effectively quantify or model human values to guide AI development? 3) what strategies have been employed to integrate human values into the AI development process? 4) how did existing studies improve human understanding and evaluation of AI alignment? 5) what are the practices for designing interfaces and interactions that facilitate human-AI collaboration? 6) How have AI been adapted to meet the needs of various human value groups? We included papers that could potentially answer any of these questions. Further, based on the scope in Section LABEL:sec:scope, we excluded papers that did not meet our inclusion criteria. This resulted in a final corpus of 411 papers, which were analyzed in detail using qualitative coding (see Appendix 8.4 for more details).
8.1.3 Qualitative Code Development.
Referring to the code development process in (lee2024design, ), we first conducted qualitative coding for each paper by identifying relevant sentences that could answer the above research questions, and entering short codes to describe them into a codebook. We iteratively coded relevant sentences from each paper through a mix of inductive and deductive approaches, which allowed flexibility to expand, modify or change the driving research questions based on our learnings as we went through the process. To ensure rigor in our coding process, two authors coded each paper. The first author independently annotated all papers after reviewing the paper abstracts and introductions. Twelve team members each annotated a subset of the paper corpus. Our corpus includes papers from different domains (e.g., HCI, NLP and ML). Therefore, we divided the authors into HCI and NLP/ML [^1] Note that NLP and ML are two different domains, we combine them together for the purposes of literature review analysis since they both work on developing and evaluating AI technologies. teams and assigned the papers accordingly based on expertise. All team members coded each of their assigned papers to answer all six questions (if applicable) introduced above.
8.1.4 Framework Development and Rigorous Coding.
After developing annotations, all authors collaborated to create the bidirectional human-AI alignment framework by integrating the annotations within each of the codes. The initial version of the framework was proposed by the author who reviewed all papers. This framework furthermore underwent iterative improvement through: 1) discussions with all team members involved in paper coding, and 2) revisions based on feedback from the project advisors. Additionally, we strengthened the framework by reviewing papers from the AI Ethics conferences (including FAccT and AIES), and related work of the collected papers that covered other domains such as psychology and social science. We further added missing codes and papers to ensure comprehensive coverage (see Appendix 8.2 for details). The final bidirectional human-AI alignment framework, with detailed topologies, is presented in Section 3. Following the framework’s finalization, we conducted another separate coding process to annotate whether each paper investigated dimensions within our framework. Two authors independently coded each paper. [^2] The joint probability of agreement for the paper annotations was 0.78. These codes were then used to perform quantitative and qualitative analyses, as presented in Section LABEL:sec:analysis.
8.2 Venues
@@
Additionally, we also consolidate the framework by reviewing the papers published in FAccT and AIES (i.e., important venues for AI Ethics research) between 2019 and 2024 and supplemented the codes, including the AI Regulatory and Policy code in Section LABEL:sec:impacts and the exemplary paper of Regulating ChatGPT hacker2023regulating ), which were not covered by the original collections. Also, we include a number of papers in the “Other” class are found by related work that are highly relevant to this topic.
8.3 Keywords
@@
8.4 Inclusion and Exclusion Criteria
@@
  • [[Human values]] we include papers that study human value definition, specification and evaluation in AI systems.
  • [[AI development techniques]] We include techniques of developing AI that aim to be more consistent with human values with interactions along all AI development stages (e.g., data collection, model construction, etc.)
  • [[AI evaluation, explanation and utilization]] we include papers that build human-AI interactive systems or conduct human studies to better evaluate, explain, and utilize AI systems.
  • [[building dataset with human interaction]] especially responsible dataset.
@@
  • [[Alignment not between human & AI]] we do not include alignment studies that are not between human and AI, such as entity alignment, cross-lingual alignment, cross-domain alignment, multi-modal alignment, token-environment alignment, etc.
  • [[AI models beyond LLMs - Modality]] we do not focus on AI models other than LLMs (e.g., 3D models, VR/AR, voice assistant, spoken assistant), our primary model modality is text. Specifically, we do not consider audio / video data; we do not consider pure computer vision modality.
  • [[No human-AI interaction]] we do not consider studies that do not involve the interaction between human and AI, such as (multi-agent) reinforcement learning. Specifically, we do not consider interactions via voices/speech, Do not consider game interaction; Do not consider interaction for Accessibility; Do not consider Mobile interaction; Not consider autonomous vehicle interaction wearable devices, or Physical interaction;
  • [[Tasks]] art and design, emotion.
  • [[No human included]]
  • [[focus on English]] primarily focus on English as the main language;
  • [[Application]] not include the NLP papers tailored for a specific traditional task, such as translation, entity recognition, sentiment analysis, knowledge graph, adversarial and defense, topic modeling, detecting AI generations, distillation, low resource, physical robots, text classification, games, image-based tasks, hate speech detection, Human Trafficking, etc.
  • [[Visualizing Embeddings]] Visualizing/interacting transformer embeddings?
  • [[Embedding-based]] explanation, evaluation, etc.
  • [[multi-agent reinforcement learning with self-play and population play]] techniques, such as self-play (SP) or population play (PP), produce agents that overfit to their training partners and do not generalize well to humans.
@@
9 Selected Paper List
10 Alignment Goals and Human Values
Table 2: The Goals of Alignment. We present the six prevailing alignment goals, associating with their Definitions (middle column), Limitations and Risks (right column). We consider Human Values as the main goal of alignment in this work referring to an extensive analysis and arguments in existing studies gabriel2020artificial ; stuart2014whitepaper
10.1 A Comprehensive Taxonomy of Human Values
This conventional theory was developed without the context of human-AI interaction, which might overlook values that need to be considered for human-AI alignment. Therefore, we used a bottom-up approach to extract all values studied in our collected alignment literature, mapped them onto the Schwartz Theory of Basic Values, and supplemented the theory with AI-related structure and content. As a result, we identified the structural relationships among human values and mapped existing literature to a fine-grained taxonomy (see Table 3). We supplemented the traditional theory’s four high-order value types (i.e., “Self-Enhancement”, “Openness to Change”, “Conservation”, “Self-Transcendence”) with a novel high-order value type, named “Desired Values for AI Tools” that encompasses two motivational value types (i.e., “Usability” and “Human-Likeness”). We further organize the relationship among these value types along two dimensions schwartz2012overview : different resources (i.e., individuals, society and interaction) and different self-intentions (i.e., self-protection against threat and self-expansion and growth). Furthermore, we elaborate the definitions of the 12 motivational value types and their exemplary values by mapping them to relevant human-AI alignment papers from our corpus in Table 3. During the process of mapping, we found: 1) value terms in empirical papers were often named differently (e.g., capability and competence), or check opposites (e.g., fairness and bias); 2) there are many values not studied in our corpus, i.e., indicated as ([Uncaptioned image]) in the Figure.
10.2 Insights into Human Values for Alignment
Table 3: The value relations and taxonomy. We consider 5 high-order value types encompassing 12 motivational value types, indicated by their sources (e.g., individuals, society and interaction).
Value Prioritization in AI Systems. Human value systems are not merely subsets of values, but ordered systems with relative priorities schwartz1994there ; schwartz2012overview . For instance, schwartz2012overview presented the definition for this phenomenon: “a value is ordered by importance relative to other values to form a system of value priorities. The relative importance of multiple values guides action….The trade-off among relevant, competing values guides attitudes and behaviors.”. Current AI alignment algorithms, often based on datasets of human preferences ouyang2022training ; song2024preference ; rafailov2024direct , may inadvertently prioritize majority values, potentially neglecting those of marginalized groups fleisig2023fairprism . Future research should address this complex interplay of values in AI systems.
Universal vs. Personalized AI Values. While certain values are universally expected from AI (e.g., capability, equity, responsibility), others may be undesirable in specific contexts ngo2022alignment (e.g., seeking power). Simultaneously, AI models should be adaptable to diverse human value systems sorensen2024roadmap . Research is needed to develop methods for identifying appropriate value sets for specific individuals or groups, and for customizing AI to align with user values while maintaining ethical principles.
Disparities in Value Expectations and Evaluation. The fundamental differences between humans and AI necessitate distinct approaches to value evaluation. For instance, assessing AI honesty may require mechanistic interpretability bereska2024mechanistic , a more rigorous standard than that applied to humans. Future studies should explore methods for evaluating and explaining AI values and calibrating human expectations accordingly.
11 Interaction Techniques for Specifying Human Values
Domain-Specific Interaction Techniques. The interaction techniques in AI-centered (NLP/ML) and Human-centered (HCI) alignment studies are often differ brath2021surveying . NLP/ML studies primarily utilize numeric and natural language-based techniques. Also, NLP/ML research explore implicit feedback to extract human hidden feedback. In contrast, HCI research encompasses a broader range of graphical and multi-modal interaction signals (e.g., sketches, location information) beyond text and images. This disparity suggests potential gaps in extracting comprehensive human behavioral information.
Figure 4: The interaction techniques for specifying values in human-AI alignment.
    Takeaways of Interaction Techniques for Alignment. 1. Some common human feedback styles used in NLP/ML are not often studied in HCI. 2. Diverse human interactive feedback in HCI are not fully used in AI development in NLP/ML.
12 Challenges in Achieving Alignment
The concept of alignment in AI research has a long history, tracing back to 1960, when AI pioneer Norbert Wiener wiener1960some described the AI alignment problem as: “If we use, to achieve our purposes, a mechanical agency with whose operation we cannot interfere effectively … we had better be quite sure that the purpose put into the machine is the purpose which we really desire.” Discussion around intelligent agents and the associated concerns relating to ethics and society have emerged since then stuurman2001software . Next, we discuss the well-known challenges encountered in achieving alignment.
Challenge 1: Outer and Inner Alignment. In the context of “intelligent agents,” until now, AI alignment research has aimed to ensure that any AI systems that would be set free to make decisions on our behalf would act appropriately and reduce unintended consequences wooldridge1999intelligent ; stuurman2001software ; russell2016artificial . At the near-term stage, aligning AI involves two main challenges: carefully specifying the purpose of the system (outer alignment i.e., providing well-specified rewards ngo2022alignment ) and ensuring that the system adopts the specification robustly (inner alignment, i.e., ensuring that every action given an agent in a particular state learns desirable internally-represented goals ngo2022alignment ). Significant efforts have been made, for inner alignment, to align AI systems to follow alignment goals of an individual or a group (e.g., instructions, preferences, values, and/or ethical principles) ouyang2022training and to evaluate the performance of alignment santurkar2023whose . However, for outer alignment, AI designers are still facing difficulties in specifying the full range of desired and undesired alignment goals of humans.
Challenge 2: Specification Gaming. To learn human alignment goals, AI designers typically provide an objective function, instructions, reward function, or feedback to the system, which is often unable to completely specify all important values and constraints that a human intended hemphill2020human . Hence, AI designers resort to easy-to-specify proxy goals such as maximizing the approval of human overseers wiki_AI_alignment , which results in “specification gaming” krakovna2020specification or “reward hacking” pan2022effects issues (i.e., AI systems can find loopholes that help them accomplish the specific objective efficiently but in unintended, possibly harmful ways). Additionally, the black-box nature of neural networks brings additional ethical and safety concerns for alignment because humans don’t know about the inner state and the actions AI leveraged to achieve the output. Consequently, AI systems might make “correct” decisions with “incorrect” reasons, which are difficult to discern. Society is already facing these issues, such as data privacy tucker2018privacy , algorithmic bias hajian2016algorithmic , self-driving car accidents bonnefon2016social , and more. As a result, these considerations necessitate considering human-AI interaction in AI alignment for specification and evaluation, ranging from addressing problems around who uses an AI system, with what goals to specify, and if the AI system perform its intended function from the user’s perspective.
Challenge 3: Scalable Oversight. From a long-term perspective, when advanced AI systems become more complex and capable (e.g., AGI morris2024levels ), it becomes increasingly difficult to align them to human values through human feedback. Evaluating complex AI behaviors applied to increasingly challenging tasks can be slow or infeasible for humans to ensure all sub-steps are aligned with their values terry2023ai . Therefore, researchers have begun to investigate how to reduce the time and effort for human supervision, and how to assist human supervisors, referred to as Scalable Oversight amodei2016concrete .
Challenge 4: Dynamic Nature. As AI systems become increasingly powerful, the alignment solutions must also adapt dynamically since human values and preferences change as well. As dautenhahn2000living posit, AI systems may be neither humane nor desirable if we do not ask questions about the long-term cognitive and social effects of social agent systems (e.g., how will agent technology affect human cognition). All these considerations call for a long-term and dynamic perspective to address human-AI alignment as an ongoing, mutual process with the collective efforts of cross-domain expertise.
Challenge 5: Existential Risk. Further, some AI researchers claim that bengio2024managing advanced AI systems will begin to seek power over their environment (e.g., humans) once deployed in real-world settings, as such behavior may not be noticed during training. For example, some language models seek power in text-based social environments by gaining money, resources, or social influence pan2023rewards . Consequently, some hypothesize that future AI, if not properly aligned with human values, could pose an existential risk to humans dafoe2016yes .
13 Author contributions